OpenStudy (anonymous):

is the following sequence arithmetic? is so identify the common difrance ..-73,-44,-15,14..... im guessing its -29

OpenStudy (texaschic101):

-44 - (-73) = -44 + 73 = 29 -15 - (-44) = -15 + 44 = 29 14 - (-15) = 14 + 15 = 29 common difference is 29....it is not negative. It IS an arithmetic sequence

OpenStudy (anonymous):

thank youuu.!!!!

OpenStudy (anonymous):

nice profile pic dude

OpenStudy (texaschic101):

glad I could help :)